Grade 11Physical Chemistry

How many moles of methane are required to produce 22 g CO2 (g) after combustion?

lily
8 Years agoGrade 11

1 Answer

Arun
8 Years ago
Dear student
 

By unitary method and stoichiometric calculation,

CH4 + 2O2 → CO2 + 2H2O

1 mol of methane → 1 mol of carbon dioxide (→ indicates “gives”)

Hence 16 g of methane → 44g of carbon dioxide

So, x (say) grams of methane → 22g of carbon dioxide

By unitary method,

x=16/44 * 22 = 8 g of methane

Converting 8 grams of methane to moles:

8/16=0.5 moles

hence your answer is ½ moles.

 

Regards

Arun (askIITians forum expert)